OpenStudy (anonymous):

Which of the following represent the zeros of f(x) = 6x^3 – 29x^2 – 6x + 5 5, 1/3, -1/2 5, -1/3, 1/2 -5, 1/3, -1/2, 5, 1/3, 1/2

OpenStudy (anonymous):

@phi

OpenStudy (phi):

I would test each choice in the original equation and see if I get 0

OpenStudy (anonymous):

To test it what do I replace? If I was testing the first one would it look like this: f(x) = 6(5)^3 – 29(1/3)^2 – 6(-1/2) + 5

OpenStudy (phi):

you replace x with the same number. so you can test +5 by replacing all x's with 5

OpenStudy (phi):

If you get 0 with x=5, that tells you choices A, B and D are possible answers. then check the next number.

OpenStudy (anonymous):

Well 6*5=30 and 30 cubed = 27000 and 29*5=145 and 145 squared = 21025 So it couldn't be those 3 could it?

OpenStudy (anonymous):

I don't think it could be -5 either though

OpenStudy (phi):

6x^3 means 6*x*x*x if x is 5, it is 6*5*5*5 or 6* 5^3 if you use a calculator

OpenStudy (anonymous):

750

OpenStudy (phi):

so you have to be careful 6x^3 – 29x^2 – 6x + 5 6*5^3 - 29* 5^2 -6*5 + 5

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

750-725-30+5=0

OpenStudy (phi):

yes, that makes sense. so you can cross of choice C now test ⅓

OpenStudy (anonymous):

I did \[6\times \frac{ 1 }{ 3 }^{3}\] and got 0.22222

OpenStudy (phi):

yes. with ⅓ I would put ⅓ into a variable in the calculator (if you know how)

OpenStudy (anonymous):

I don't...

OpenStudy (phi):

ok, then keep going... -29*(⅓)^2

OpenStudy (anonymous):

-3.22222...

OpenStudy (phi):

if you add the first term 0.22222 you get -3 so far. now – 6x

OpenStudy (anonymous):

I put the fraction in parenthesis and got -3.625

OpenStudy (phi):

I don't know how you got -3.625 so far you found 6 *(⅓)^3 = 0.22222 -29*(⅓)^2 = -3.22222 which add up to -3 now -6x+5

OpenStudy (anonymous):

Wouldn't it add up to -3.44444

OpenStudy (phi):

+0.2222 - 3.2222 maybe it is easier to see if you look at a similar problem +0.5 - 1.5 (plus ½ take away 1 and ½ , leaves -½)

OpenStudy (anonymous):

-6*1/3=-2 -3+-2=-5 5+-5=0

OpenStudy (phi):

the first line is ok -6*⅓ is -2 now add +5 +5 -2 is +3 all together, 6 *(⅓)^3 = 0.22222 -29*(⅓)^2 = -3.22222 -6x = -2 +5 = +5 now add up all of those numbers.

OpenStudy (anonymous):

6*-1/2^3=-0.75 29*-1/2^2=-7.25 6*-1/2=-3+5 -0.75-7.25-3+5=-6

OpenStudy (anonymous):

I added those already above

OpenStudy (phi):

I think you have to be careful of the signs. notice in the equation it is -6*x if x is -½ , this is -6*-½ = +3 (not -3)

OpenStudy (anonymous):

Ohhh

OpenStudy (anonymous):

So it would be 0 :D

OpenStudy (anonymous):

Thanks Phi!

OpenStudy (anonymous):

I have another question though...

OpenStudy (anonymous):

It's unrelated to this one

OpenStudy (anonymous):

Which polynomial identity will prove that 37 = 64 − 27

OpenStudy (phi):

yes, and your second line 29*-1/2^2=-7.25 is the correct number, but you should have said -29* (-½)^2 the -½ * -½ = +¼ , and -29* ¼= -7.25

OpenStudy (anonymous):

My answer choices are: difference of squares difference of cubes sum of cubes square of a binomial I would guess it's sum of cubes because it's adding but there are no cubes...

OpenStudy (phi):

4*4*4 and 3*3*3

OpenStudy (anonymous):

What do you mean?

OpenStudy (phi):

what do you get if you multiply out 4*4*4 or 3*3*3 ?

OpenStudy (anonymous):

Maybe it's not sum of cubes, it's not adding. Idk what i was thinking

OpenStudy (anonymous):

64 and 27 which are cubes

OpenStudy (anonymous):

and its subtracting

OpenStudy (anonymous):

is it difference of cubes?

OpenStudy (phi):

that would be my guess.

OpenStudy (anonymous):

So would 49 − 4 = 45 be difference of cubes?

OpenStudy (anonymous):

49 and 4 are cubes and its subtracting

OpenStudy (phi):

looks like it: 7^2 - 2^2 - 45

OpenStudy (phi):

=45 that looks like a difference of squares.

OpenStudy (anonymous):

You were right, thank you :)